# CER-HV: A CER-Based Human-in-the-Loop Framework for Cleaning Datasets Applied to Arabic-Script HTR

### v2

@@ 3.1 Problem Formulation
In CTC based models such as the CRNN, the network outputs a sequence of frame level probability distributions over the extended alphabet $\Sigma^{\prime}=\Sigma\cup\{\text{blank}\}$. The blank symbol allows the model to handle variable length alignments by representing either no character or repeated characters. Let $T$ be the length of the output sequence produced by the encoder. Each alignment path $\pi=(\pi_{1},\dots,\pi_{T})$ belongs to $(\Sigma^{\prime})^{T}$ and collapses to a final transcription through a function $B$ that removes blanks and merges repeated characters.
The probability of $y$ is defined as the sum over all valid alignment paths:
@@ 3.2 HTR Model: CRNN Architecture
Our architecture builds upon the original CRNN framework proposed by Shi et al. [30], which integrates convolutional, recurrent, and transcription components into a unified end-to-end model. We adopt the refined implementation introduced in the Best Practices framework [25], which enhances the CRNN design with deeper residual layers, normalization, and improved training strategies. As illustrated in Figure 3, the model follows an encoder–decoder structure, consisting of a CNN encoder for visual feature extraction and an RNN decoder for sequence modeling and transcription.
@@
The CNN encoder extracts hierarchical visual features from the input text line image and converts the resulting two-dimensional feature maps into a one-dimensional sequence along the image width using column-wise max pooling. This preserves the horizontal reading order while providing a compact representation suitable for recurrent sequence modeling. The recurrent decoder then models contextual dependencies and produces per-frame character probabilities, which are decoded using CTC.
While most Arabic-script HTR systems adopt CRNN structures combining convolutional feature extraction, recurrent sequence modeling, and CTC decoding, implementation details significantly affect performance. Earlier approaches such as Shi et al. [30] and Puigcerver [24] employed shallow VGG-style convolutional backbones that limited fine-grained stroke feature capture. Following the Best Practices framework [25], our implementation incorporates three critical modifications: (1) a deeper residual CNN encoder with batch normalization for improved feature learning, (2) column-wise max pooling instead of concatenation to reduce feature dimensionality while maintaining vertical invariance, and (3) an auxiliary CTC shortcut branch during training to enhance recurrent layer convergence.
Concretely, the CNN encoder begins with a $\times7$ convolution (32 channels, stride $\times2$), followed by ResNet blocks (2 with 64 channels, 3 with 128, and 2 with 256), interleaved with batch normalization and dropout. Progressive $\times2$ max-pooling between block groups downsamples the spatial resolution, producing feature maps of size (H/8) $\times$ (W/8) $\times$ 256. Column-wise max pooling then flattens this tensor into a sequence of (W/8) feature vectors (256 dimensions each).
The recurrent head and decoding stage use three stacked BiLSTM layers (256 hidden units each) to process the feature sequence. A fully connected classifier projects each BiLSTM output onto the character space, followed by softmax normalization to obtain per-frame probability distributions. During inference, greedy CTC decoding selects the most probable character at each time step, merges duplicates, and removes blanks to generate the final transcription.
@@ 3.3 CER-HV Framework: CER Based Ranking and HITL Verification
The second component of the methodology focuses on improving data quality through automated label error detection and human verification. The CRNN recognizer depends critically on the quality of the training data. Potential label noise such as incorrect transcriptions, image segmentation errors, or irrelevant samples can significantly degrade accuracy [22]. To address this, we introduce CER-HV, a two-stage framework that combines automated label noise detection based on CER scoring with human verification. Figure 2 shows the overall flow.
@@
Second, rather than using the cyclical multi-round training procedure of O2U-Net to avoid overfitting, we employ an early-stopping criterion. Prior studies have shown that early stopping based on validation metrics effectively detects convergence and prevents overfitting [32]. We identify a single, objectively defined convergence epoch $t_{\mathrm{conv}}$ at which training and validation CER values converge and further improvement becomes negligible. This choice simplifies the procedure, removes the need for hyperparameter tuning, and aligns better with sequence-prediction tasks such as HTR.
Once the CRNN reaches the convergence epoch, each training sample is assigned a score
@@
    Transcription error– the ground truth text does not match the handwritten content in the image.
@@
    Segmentation error– multiple text lines are included within a single image, or the lines are truncated/incomplete.
@@
    Orientation error– significantly rotated text lines that contradict the expected reading direction.
@@
    Script mismatch– the content is written in a script or language different from the target alphabet (e.g., Latin or numeric text appearing in Arabic-script datasets).
@@
    Irrelevant or non-text samples– stamps, signatures or other non-textual document elements.
@@
    Valid but hard– the sample is correct, but the sample is visually challenging to correctly automatically transcribe.
@@ 4.1 Datasets
Dataset Language Train/Val/Test Avg. Width–Height Max Char Len
@@
Table 1: Dataset descriptions including train/val/test split sizes, followed by average line image dimensions and character length ranges.
@@
PHTI (Pashto): The Pashto Handwritten Text Imagebase (PHTI) is a dataset containing 36,082 handwritten text-line images representing real-world Pashto text, along with ground truth annotations in UTF-8 encoding. The dataset encompasses diverse genres including short stories, historical narratives, poetry, and religious content, collected from a variety of modern writers.
@@
Because several pages contain highly similar or near-duplicate content, we performed a similarity analysis to prevent data leakage between training and evaluation splits. Duplicate detection revealed that 2 pages were exact duplicates, and among the 9,730 possible page pairs, 1,155 pairs (11.9%) exhibited more than 85% textual similarity. Only 37 pages showed no significant overlap with any other page. The validation and test sets were selected exclusively from these 37 non-overlapping pages to ensure that no evaluation sample is textually related to the training data.
The final split contains 110 pages for training (1,473 line images), 15 pages for validation (160 lines), and 15 pages for testing (155 lines), corresponding to an 80%,10%,10% division applied at the page level to ensure a leakage-free evaluation protocol for Persian handwritten text recognition. The cleaned line images and the final train, validation, and test splits will be released publicly [^2] PHTD cleaned dataset: https://huggingface.co/datasets/sana-ngu/PHTD.
@@
Critically, to prevent data leakage between training and evaluation splits, we performed a similarity check and ensured that no samples were shared between the validation and test sets. The baseline Arabic HTR models used for evaluation by the dataset creators were trained on Middle Eastern and South Asian manuscripts, and yielded high CERs ranging from 65–84% when applied to this corpus. We prepared and publicly released the extracted line images and data splits used in this work as an open-source resource to ensure reproducibility for future research.
@@ 4.2 Training Setup and Evaluation
For each dataset, image preprocessing followed a standardized resizing and padding procedure to reduce shape distortion and ensure uniform input dimensions. Specifically, all line images were resized to the rounded average height and width (in whole pixels) of their respective training set, as reported in Table 1. If an image exceeded these target dimensions, it was proportionally rescaled; otherwise, it was centered and padded with median pixel values. This approach preserved the aspect ratio while providing consistent spatial context across samples, which we observed stabilized training and improved recognition performance on longer text lines.
@@
Recognition performance was assessed using the Character Error Rate (CER), the standard metric in HTR. CER is computed as the normalized Levenshtein (edit) distance between the predicted and ground-truth transcriptions, reflecting the proportion of character insertions, deletions, and substitutions required to match the prediction to the reference. Lower CER indicates higher transcription accuracy and provides a fine-grained measure of model performance at the character level.
@@ 5 Results
This section presents the experimental results of our study, focusing on baseline CRNN performance and the impact of dataset label quality. We also assess the effectiveness of the proposed CER-HV framework.
We begin by comparing our CRNN results with prior work across all datasets using the original, unaltered data to establish quantitative CER baselines. We then apply the CER-HV framework to analyze label noise across the training, validation, and test splits. Representative qualitative examples of common labeling issues are presented to illustrate the identified error categories. Finally, we quantify the impact of correcting these errors by evaluating how dataset cleaning influences recognition performance.
@@ 5.1 Quantitative Benchmarking of the CRNN-Based HTR System
For PHTI, our CRNN achieves a substantial improvement over prior non-transformer models, reducing CER from 20.7% to 8.26%, representing one of the largest performance gains across all evaluated datasets. For PHTD, where publicly reported benchmarks are limited, our model establishes a new baseline of 11.3% CER. On NUST-UHWR, our model reaches 6.66% CER, improving upon earlier convolutional and recurrent baselines [34, 26, 12], although Transformer-based architectures [19, 12] still achieve superior performance due to attention mechanisms and synthetic pretraining. For Ajami, the model attains 10.66% CER, far outperforming previously reported results (64-84%), which were obtained through zero-shot evaluation using pretrained OPEN-ITI models [38].
@@
Conv-Transformer [26] No 6.4
Ajami (Fulfulde/Hausa) [38] Pre-trained models from OPEN ITI [38] No 64–84
@@ 5.2 Label Error Analysis
Dataset
Split
Transcription Error
Segmentation Error
Orientation Error
Script Mismatch
Irrelevant Content
Total
@@
Table 4: Distribution of detected label errors by type across training, validation, and test splits.
@@ 5.4 Quantitative Impact of Label Errors
Dataset Training Set Evaluation Set Validation Split Test Split
@@
These results show that even limited amounts of label noise can substantially distort model assessment, emphasizing the need for systematic validation of evaluation sets to ensure reliable and fair benchmarking in HTR. While evaluation-set cleaning is critical for obtaining trustworthy performance estimates, training-set cleaning provides measurable benefits primarily when noise levels are sufficiently high and dominated by structurally disruptive errors that interfere with sequence learning. In settings where training noise is sparse or mild, retraining on cleaned data yields diminishing returns.
@@ 6 Conclusion and Future Work
The effectiveness of the human-in-the-loop component in CER-HV is closely tied to the characteristics of handwritten text recognition. In CTC-based models, loss values are an indirect and often unreliable indicator of label correctness, whereas CER directly reflects transcription quality and is immediately interpretable by human reviewers. Because line-level HTR datasets typically contain thousands rather than millions of samples, restricting inspection to a small subset of high-CER samples makes targeted verification feasible and low cost. This lightweight process allows obvious structural errors—such as segmentation failures, orientation mistakes, script mismatch, or non-text content—to be identified efficiently without exhaustive manual review.
Beyond its use in analyzing existing benchmarks, CER-HV can serve as a practical dataset validation tool. Applied during dataset creation or prior to benchmarking, it enables researchers to identify likely mislabeled or out-of-scope samples and verify only a small number of high-risk cases. Our findings further highlight that data quality represents a complementary dimension to architectural innovation: when annotation issues dominate, increasing model complexity does not reliably improve performance and may instead reinforce incorrect supervision.
A limitation of the present study suggests directions for future work. First, cleaning was implemented through exclusion of noisy samples rather than correction. While this strategy enables a controlled analysis of how label noise affects measured recognition accuracy, it does not exploit the full information contained in the identified samples. Different error types would require different remedies, such as pseudo-labeling for transcription errors, dedicated segmentation models for segmentation errors, or automatic rotation strategies for orientation errors. The released CER-HV error annotations and trained recognition models provide a practical foundation for active learning and semi-automated correction methods in future research.

### v3

# CER-HV: A Human-in-the-Loop Framework for Cleaning Datasets Applied to Arabic-Script HTR

@@ 3.1 Problem Formulation
In CTC based models such as the CRNN, the network outputs a sequence of frame level probability distributions over the extended alphabet $\Sigma^{\prime}=\Sigma\cup\{\text{blank}\}$. The blank symbol allows the model to handle variable length alignments by representing either no character or repeated characters. Let $T$ be the length of the output sequence produced by the encoder. Each alignment path $\pi=(\pi_{1},\dots,\pi_{T})$ belongs to $(\Sigma^{\prime})^{T}$ and collapses to a final transcription through a function $B$ that removes blanks and merges repeated characters. The probability of $y$ is defined as the sum over all valid alignment paths:
@@ 3.2 HTR Model: CRNN Architecture
Our architecture builds upon the original CRNN framework proposed by Shi et al. [30], which integrates convolutional, recurrent, and transcription components into a unified end-to-end model. We adopt the refined implementation introduced in the Best Practices framework [25], which enhances the CRNN design with deeper residual layers, normalization, and improved training strategies. The overall architecture is illustrated in Figure 3. It consists of a convolutional feature extractor followed by a recurrent sequence modeling module trained with CTC
@@
The CNN feature extractor learns hierarchical visual representations from the input text line image and converts the resulting two-dimensional feature maps into a one-dimensional sequence along the image width using column-wise max pooling. This preserves the horizontal reading order while providing a compact representation suitable for sequence processing. The recurrent sequence modeling module then captures contextual dependencies and produces per-frame character probabilities, which are decoded using CTC.
While most Arabic-script HTR systems adopt CRNN structures combining convolutional feature extraction, recurrent sequence modeling, and CTC decoding, implementation details significantly affect performance. Earlier approaches such as Shi et al. [30] and Puigcerver [24] employed shallow VGG-style convolutional backbones that limited fine-grained stroke feature capture. Following the Best Practices framework [25], our implementation incorporates three critical modifications: (1) a deeper residual feature extractor with batch normalization for improved feature learning, (2) column-wise max pooling instead of concatenation to reduce feature dimensionality while maintaining vertical invariance, and (3) an auxiliary CTC shortcut branch during training to enhance recurrent layer convergence.
Concretely, the CNN feature extractor begins with a $\times7$ convolution (32 channels, stride $\times2$), followed by ResNet blocks (2 with 64 channels, 4 with 128, and 4 with 256), interleaved with batch normalization and dropout. Progressive $\times2$ max-pooling between block groups downsamples the spatial resolution, producing feature maps of size (H/8) $\times$ (W/8) $\times$ 256. Column-wise max pooling then flattens this tensor into a sequence of (W/8) feature vectors (256 dimensions each).
The recurrent sequence modeling module consists of three stacked BiLSTM layers (256 hidden units each), followed by a linear projection onto the character space. During training, CTC is applied to the resulting character logits. During inference, greedy CTC decoding selects the most probable character at each time step, merges duplicates, and removes blanks to generate the final transcription.
@@ 3.3 CER-HV Framework: CER Based Ranking and HITL Verification
The second component of the methodology focuses on improving data quality through automated label error detection and human verification. The CRNN recognizer depends critically on the quality of the training data. Potential label noise can significantly degrade accuracy [22]. To address this, we introduce CER-HV, a two-stage framework that combines automated label noise detection based on CER scoring with human verification. Figure 2 shows the overall flow.
@@
Second, rather than using the cyclical multi-round training procedure of O2U-Net to avoid overfitting, we employ an early-stopping criterion. Prior studies have shown that early stopping based on validation metrics effectively detects convergence and prevents overfitting [32]. We identify a single, objectively defined convergence epoch $t_{\mathrm{conv}}$ at which training and validation CER values converge and further improvement becomes negligible. This choice simplifies the procedure, removes the need for hyperparameter tuning, and aligns better with sequence-prediction tasks such as HTR. Once the CRNN reaches the convergence epoch, each training sample is assigned a score
@@
    Transcription error (TE)– the ground truth text does not match the handwritten content in the image.
@@
    Segmentation error (SE)– multiple text lines are included within a single image, or the lines are truncated/incomplete.
@@
    Orientation error (OE)– significantly rotated text lines that contradict the expected reading direction.
@@
    Script mismatch (SM)– the content is written in a script or language different from the target alphabet (e.g., Latin or numeric text in Arabic-script datasets).
@@
    Irrelevant or non-text content (IC)– stamps, signatures or other non-textual document elements.
@@
    Valid but hard (VA)– the sample is correct, but the sample is visually challenging to correctly automatically transcribe.
@@ 4.1 Datasets
Dataset Language Train/Val/Test Avg. W–H Max Len
@@
Table 1: Dataset descriptions including train/val/test split sizes, followed by average line image dimensions (Width–Height) and character maximum length.
@@
PHTI (Pashto): The Pashto Handwritten Text Imagebase (PHTI) is a dataset containing 36,082 handwritten text-line images representing real-world Pashto text, along with ground truth annotations. The dataset encompasses diverse genres including short stories, historical narratives, poetry, and religious content, collected from a variety of modern writers.
@@
Because several pages contain highly similar or near-duplicate content, we performed a similarity analysis to prevent data leakage between training and evaluation splits. Duplicate detection revealed that 2 pages were exact duplicates, and among the 9,730 possible page pairs, 1,155 pairs (11.9%) exhibited more than 85% textual similarity. Only 37 pages showed no significant overlap with any other page. The validation and test sets were selected exclusively from these 37 non-overlapping pages. The cleaned line images and the final train, validation, and test splits will be released publicly [^2] PHTD cleaned dataset: https://huggingface.co/datasets/sana-ngu/PHTD.
@@
To prevent data leakage between training and evaluation splits, we performed a similarity check and ensured that no samples were shared between the validation and test sets. The baseline Arabic HTR models used for evaluation by the dataset creators were trained on Middle Eastern and South Asian manuscripts, and yielded high CERs ranging from 65–84% when applied to this corpus. We prepared and publicly released the extracted line images and data splits used in this work as an open-source resource to ensure reproducibility for future research.
@@ 4.2 Training Setup and Evaluation
For each dataset, image preprocessing followed a standardized resizing and padding procedure to reduce shape distortion and preserve aspect ratio. All line images were resized to the rounded average height and width (in whole pixels) reported in Table 1. Images exceeding these dimensions were proportionally rescaled; otherwise, they were centered and padded with median pixel values.
@@
Recognition performance was assessed using the Character Error Rate (CER), the standard metric in HTR. CER is computed as the normalized Levenshtein (edit) distance between the predicted and ground-truth transcriptions, reflecting the proportion of character insertions, deletions, and substitutions required to match the prediction to the reference. Lower CER indicates higher transcription accuracy.
@@ 5 Results
This section presents the experimental results of our study, we begin by comparing our CRNN results with prior work across all datasets using the original, unaltered data to establish quantitative CER baselines. We then apply the CER-HV framework to analyze label noise across the training, validation, and test splits. Representative qualitative examples of common labeling issues are presented to illustrate the identified error categories. Finally, we quantify the impact of correcting these errors by evaluating how dataset cleaning influences recognition performance.
@@ 5.1 Quantitative Benchmarking of the CRNN-Based HTR System
For PHTI, our CRNN achieves a substantial improvement over prior non-transformer models, reducing CER from 20.7% to 8.26%, representing one of the largest performance gains across all evaluated datasets. For PHTD, where publicly reported benchmarks are limited, our model establishes a new baseline of 11.3% CER. On NUST-UHWR, our model reaches 6.66% CER, improving upon earlier convolutional and recurrent baselines [34, 12], although Transformer-based architectures [19, 12] still achieve superior performance due to attention mechanisms and synthetic pretraining. For Ajami, the model attains 10.66% CER, far outperforming previously reported results (64-84%), which were obtained through zero-shot evaluation using pretrained OPEN-ITI models [38].
@@
Ajami (Fulfulde/Hausa) [38] OPEN-ITI[38] No 64–84
@@ 5.2 Label Error Analysis
Dataset Split TE SE OE SM IC Total
@@
Table 4: Distribution of detected label errors by type across training, validation, and test splits. TE: Transcription Error; SE: Segmentation Error; OE: Orientation Error; SM: Script Mismatch; IC: Irrelevant Content.
@@ 5.4 Quantitative Impact of Label Errors
Dataset Training Set Eval. Set Validation Split Test Split
@@
These results show that even limited amounts of label noise can substantially distort model assessment, emphasizing the need for systematic validation of evaluation sets to ensure reliable and fair benchmarking in HTR.
@@ 6 Conclusion and Future Work
The effectiveness of the human-in-the-loop component in CER-HV is closely tied to the characteristics of HTR. In CTC-based models, loss values are an indirect and often unreliable indicator of label correctness, whereas CER directly reflects transcription quality and is immediately interpretable by human reviewers. Because line-level HTR datasets typically contain thousands rather than millions of samples, restricting inspection to a small subset of high-CER samples makes targeted verification feasible and low cost. This lightweight process allows obvious structural errors—such as segmentation failures, orientation mistakes, script mismatch, or non-text content—to be identified efficiently without exhaustive manual review.
Beyond its use in analyzing existing benchmarks, CER-HV can serve as a practical dataset validation tool. Applied during dataset creation, it enables researchers to identify likely mislabeled or out-of-scope samples and verify only a small number of high-risk cases. Our findings further highlight that data quality represents a complementary dimension to architectural innovation: when annotation issues dominate, increasing model complexity does not reliably improve performance and may instead reinforce incorrect supervision.
A limitation of the present study suggests directions for future work. First, cleaning was implemented through exclusion of noisy samples rather than correction. While this strategy enables a controlled analysis of how label noise affects measured recognition accuracy, it does not exploit the full information contained in the identified samples. Different error types would require different remedies, such as pseudo-labeling for transcription errors, dedicated segmentation models for segmentation errors, or automatic rotation strategies for orientation errors. The released CER-HV error annotations provide a practical foundation for active learning and semi-automated correction methods in future research.
@@
7 CRediT authorship contribution statement
Sana Al-azzawi: Conceptualization, Methodology, Software , Visualization, Data curation, Formal analysis, Writing- Original draft preparation. Elisa Barney: Validation, Supervision, Conceptualization, Writing- Reviewing and Editing. Marcus liwicki: Supervision, Project administration, Writing- Reviewing and Editing.
@@ 8 Acknowledgements
9 Data Availability
The datasets used in this study are publicly available